# Can industrial overcapacity enable seasonal flexibility in electricity use? A case study of aluminum smelting in China

### v1

@@ Abstract
In many countries, declining demand in energy-intensive industries (EIIs) such as cement, steel, and aluminum is leading to industrial overcapacity. Although industrial overcapacity is traditionally envisioned as problematic and resource-wasteful, it could unlock EIIs’ flexibility in electricity use. Here, using China’s aluminum smelting industry as a case, we evaluate the system-level cost-benefit of retaining EII overcapacity for flexible electricity use in decarbonized energy systems. We find that overcapacity can enable aluminum smelters to adopt a seasonal operation paradigm, ceasing production during winter load peaks exacerbated by heating electrification and renewable seasonality. In the 2050-net-zero scenario, the seasonal operation paradigm can reduce China’s electricity system investment and operational costs by 15-72 billion CNY/year (or 8-34% of the aluminum smelting industry’s product value), sufficient to offset the costs of maintaining overcapacity and product storage. It also reduces workforce fluctuations across the aluminum smelting and thermal power generation sectors by up to 62%, potentially mitigating socio-economic disruptions from industrial restructuring and energy transition.
@@ 1 Introduction
Industrial restructuring — the transition of economies from heavy industry to higher-added-value sectors [tan2024toward] — is leading to declining demand for industrial products and overcapacity in energy-intensive industries in many countries. China’s steel and cement industries, for instance, have averaged 20-30% overcapacity in the last decade, resulting in declining profitability and asset utilization [xie2023impact]. China’s aluminum smelters will confront substantial excess capacity in the near future, as China’s domestic aluminum demand is increasingly met by recycled aluminum (potentially fulfilling 70% of future demand [li_when_2020]). Traditionally, this structural overcapacity is viewed as problematic, causing resource waste and economic risks [yu2020environmental]. Consequently, governments typically strive to eliminate overcapacity through policy measures [guo2022uncovering], such as China’s Supply-Side Structural Reform (SSSR), which was initiated around 2016 to tackle massive industrial overcapacity.
@@
On the other hand, excess capacity also unlocks operational flexibility for energy-intensive industries, as it frees them from operating at full capacity at all times to meet demand. In decarbonized power systems reliant on variable renewable energy, the demand-side flexibility of energy-intensive industries may become more valuable for balancing electricity supply and demand [cole_quantifying_2021, jenkins_long-duration_2021, wei_generation_2025]. By shifting electricity consumption to periods of high renewable energy availability, energy-intensive industries can reduce both the energy system operational costs and carbon emissions [golmohamadi_demand-side_2022]. Furthermore, unlike “fast-burst” demand-side resources such as electric vehicles and air conditioning [sepulveda_role_2018, van_heerden_demand-side_2025], industrial facilities with overcapacity and material storage could potentially provide long-duration flexibility by scheduling their production over longer time horizons. In decarbonized energy systems with increasing variability in electricity prices [Bistline2021VariabilityID], this flexibility is also essential for lowering the energy input costs of energy-intensive industries like aluminum smelting, where electricity costs exceed 30% of operational expenses [CHALCO2020AnnualReport].
In this context, determining whether industrial overcapacity should be decommissioned or retained requires a system-level cost-benefit analysis of the two competing mechanisms: decommissioning excess capacity to reduce maintenance costs versus retaining it to reduce energy system costs and emissions. To quantitatively assess this trade-off, we select China’s aluminum smelting industry as a representative case. China is likely to face over 50% overcapacity in the aluminum smelting industry by 2050 [li_when_2020], due to a shift towards secondary production using aluminum scrap [song_chinas_2023, eheliyagoda2022role]. Furthermore, the energy input for the aluminum smelting industry is almost entirely from electricity, accounting for 7% of China’s electricity demand. The aluminum smelting industry also has technical capability for flexible operation [todd2008providing]. These make it an ideal example of an industry for examining the benefits of flexibly operating in the context of a decarbonized electricity system.
Figure 1: Conceptual framework for overcapacity-enabled seasonal flexibility in energy-intensive industries. The left panel illustrates the spring to autumn time period when renewable energy generation is abundant, with aluminum smelters operating at full capacity, producing aluminum for both meeting immediate demand and storing in inventory. The right panel illustrates winter scenarios where renewable generation is limited and heating electrification exacerbates peak demand, compelling coal and natural gas power plants to operate. During this period, aluminum smelters shut down and meet demand by drawing from product inventory. This seasonal operational paradigm leverages excess capacity to create complementary patterns between industrial production and energy system dynamics, enabling cost-effective decarbonization.
Existing studies have demonstrated the rapid power adjusting capabilities of aluminum electrolytic cells, validating the technical feasibility for flexible operation [todd2008providing, wong_studies_2023]. While existing studies analyze plant-level thermal balance and restart processes of aluminum electrolytic cells [tabereaux_loss_2016, driscoll_economics_2016], a comprehensive assessment of the system-level value of this flexibility remains absent, particularly its combined impact on energy system and aluminum smelter operation. Understanding these integrated impacts and their key determinants is essential for charting a sustainable path for the aluminum industry. However, the computational complexity of modeling and optimizing the intricate coupling between two already large and complex systems—the electricity system and aluminum smelters—has been a significant barrier to such analysis.
@@
In this work, we quantitatively evaluate the system-level flexibility value of retaining aluminum smelting overcapacity in China’s decarbonization pathway by embedding a refined aluminum smelter operation model [shen_improved_2025] within a provincial-resolved power system capacity expansion model to co-optimize the hourly operations of both systems. The problem is made computationally tractable by the dimensionality reduction framework developed by Lyu et al. [lyu_data-driven_2025] and a tailored iterative solution algorithm. We find that in net-zero electricity systems, overcapacity enables aluminum smelters to shift operations away from periods of winter electricity supply scarcity, which are driven by low renewable availability and heating electrification (Figure 1). This flexibility lowers overall electricity system costs by reducing peak loads, which in turn decreases the renewable capacity required to meet those peaks and lowers utilization of coal, natural gas, and nuclear generation. In a 2050 net-zero scenario, retaining 30% aluminum-smelting overcapacity reduces electricity system costs by 15-72 billion CNY/year (or 8-34% of the aluminum smelting industry’s product value), as compared to a no-overcapacity case. We also find that, in such a scenario, aluminum production costs are lower by over 1,500 CNY/tonne, due to aluminum smelters consuming cheaper electricity. Furthermore, because thermal power plants may also operate seasonally in decarbonized systems, complementary seasonal operation in the aluminum smelting industry could help reduce workforce fluctuations across the aluminum smelting and power generation sectors by 62%, potentially helping to stabilize employment.
@@
Although we are using aluminum smelting as a case, our findings can be generalized to other energy-intensive industries, such as steel and cement industries, where seasonal operation could potentially bring significant energy input cost savings to offset the costs of maintaining overcapacity and product storage. More broadly, our work suggests that industrial overcapacity should be reframed from a problem to an opportunity to reduce energy system costs, lower industrial production costs, and mitigate socio-economic disruptions from energy decarbonization.
@@ 2 Results
Smelter operational flexibility
Case Min. power consumption (per unit of nameplate value) Restart costs (CNY/MW)
Low $0.9$ 760,000
Mid $0.7$ 110,000
High $0.5$ 22,000
Unconstrained $0.0$ 0
@@
Technology costs (per unit of core-scenario costs)
Case Wind/solar CAPEX Battery CAPEX Long-duration storage CAPEX Methanation /direct air capture costs
Low 0.8 0.8 0.8 0.8
Mid 1 1 1 1
High 1 1.5 1.5 1.5
The normalized costs for restarting potlines of 110,000 CNY/MW [Hu2009] and 760,000 CNY/MW [noauthor_high_2020, noauthor_alcoa_2021] are based on historical events. The restart-up costs are measured from the electricity input side. As for the Mid-demand scenario, we set the per capita aluminum stock demand saturation at the level of U.S. demand (approximately 600 kg per capita). The historical scrap aluminum recovery rate in 2024 is 70% [ChinaNonferrous2025], which is used in the Mid-demand scenario. The electricity system cost data is primarily from the PyPSA-China project, originally established and open-sourced by Zhou et al. [zhou_multienergy_2024]. Long-duration storage includes H2 and water storage.
@@ 2.1 Inevitable overcapacity in aluminum smelting in the recycle era
Figure 2: China’s projected aluminum demand and smelting overcapacity under different scenarios. The top solid line represents total aluminum demand, the green area shows recycled aluminum production (RA), and the yellow area indicates primary aluminum production (PA) up to 2050. China’s aluminum smelting capacity at the end of 2024 (45 million tonnes (Mt) per year) is represented by the red dashed line. The area between the red dashed line (smelter capacity at 2024 level) and the yellow area represents the excess capacity of aluminum smelters unless decommissioned. The overcapacity rate is represented by the purple line, and is given by one minus the utilization rate (1 - demand/capacity). Different scenarios are constructed based on various assumptions regarding economic development and aluminum product import/export levels, with the variable settings outlined in Table 1. Even under the high-demand scenario, China’s aluminum industry would retain 46% excess capacity — and as much as 86% under the low-demand scenario — posing a significant challenge to the sector.
This study employs a scenario analysis approach to explore the impact of uncertain parameters on the value of overcapacity in providing flexibility in the energy system. We evaluate three key dimensions: primary aluminum demand, smelter operational flexibility, and electricity system technology costs (Table 1). We assume that annual primary aluminum demand equals total aluminum demand minus demand met by recycled aluminum, which is prioritized due to its significantly lower costs and lower carbon emissions compared to smelting [li_analysis_2021, peng_life_2022], but also constrained by scrap availability [li_when_2020]. Detailed scenario settings and considerations are provided in the Methods section and Supplementary Notes.
@@
Our analysis shows that over the next one to two decades, China’s primary aluminum demand will rapidly decline and remain low as more aluminum is produced from scrap, leading to unprecedented overcapacity in the aluminum smelting industry. As shown in Figure 2, the proportion of recycled aluminum production in total aluminum demand will increase from 25% in 2024 [ChinaNonferrous2025], to the United States level of exceeding 60% by 2050 [AluminumAssociation2024KPI]. While this transition entails substantial reductions in the industry’s energy consumption and carbon emissions, due to the lower energy intensity of producing aluminum from scrap [li_analysis_2021], it would result in over 60% overcapacity for existing aluminum smelters. The rapid increase in China’s recycled aluminum output stems from a significant increase in available scrap aluminum, which is a delayed effect of China’s climbing aluminum production over the past two decades, as aluminum products typically have a lifespan of 10-20 years [hatayama2012evolution].
From a modeling perspective, the projections for available scrap aluminum and, thus, recycled aluminum production are a calculated result of China’s historical aluminum production and are not significantly affected by model assumptions. For example, the model uses aluminum production data from 1990 to 2023 to calculate scrap aluminum availability, resulting in approximately 15 million tonnes (Mt) of available scrap aluminum in 2024. Then, the projected recycled aluminum production in 2024 is 10.5 Mt, which is the available scrap multiplied by the 70% scrap aluminum recovery rate. The number is very close to the historical production of 10.6 Mt [ChinaNonferrous2025].
In contrast, total aluminum demand is more susceptible to uncertain factors such as China’s economic development level, as well as the import and export amount of aluminum products. Nevertheless, even in the most optimistic scenario for total aluminum demand, the nationwide annual primary aluminum demand is projected to fall to 24 Mt by 2050, which is less than 54% of the 2024 level.
2.2 Electricity system cost reduction as a result of overcapacity-enabled flexibility
@@
Figure 3: Reduction in electricity system costs due to overcapacity-enabled flexibility in aluminum smelting, compared to the no-overcapacity case. Colors represent different sources of changes in system cost: renewable investment costs (green) and non-renewable operational costs (orange) account for the largest system cost reductions. The reduction in renewable investment costs reflects improved renewable energy utilization efficiency through concentrating smelting during periods of high renewable generation. The reduction in non-renewable operational costs reflects a decreased reliance on firm generation resources (coal, gas, biomass, and nuclear) to ensure power supply reliability. Cost reductions range from 15 to 72 billion CNY (8-34% of the aluminum smelting industry’s product value in 2050) across scenarios with different aluminum demand, smelter flexibility, and technology costs. The only increased investment is in heating electrification. By mitigating winter peak loads, flexible smelter operation reduces the need for costly clean firm generation and long-duration storage, thereby lowering the grid integration barriers for heating electrification.
@@
Figure 3 shows the reduction in electricity system costs in a 2050 net-zero emissions scenario when the operation of aluminum smelters with overcapacity is co-optimized with the electric power system, compared to a no-overcapacity operation scenario, which represents the value of flexibility in smelting electricity use. Here, the power system cost includes operational costs and the annualized capital investment for new generation, storage, and transmission capacity. The costs of the aluminum smelters are not included in the electricity system costs, but are discussed later in the paper. The results show that the overcapacity of aluminum smelters adds a flexibility value of 15 to 72 billion CNY to the power system, which is equivalent to 8-34% of the aluminum smelting industry’s product value in 2050. National aggregated results are shown here, while a provincial resolution for smelter and electricity system modeling is implemented (see the Methods section and Supplementary Notes 3.2 and 4.2 for details).
Across all scenarios, the cost reductions are primarily driven by lower investment costs in renewable energy and lower operational costs for coal, natural gas, and nuclear energy (mainly fuel costs). Two mechanisms can explain the dominance of these cost reduction sources. First, wind and solar power exhibit high variability, resulting in low utilization rates without flexibility mechanisms such as storage or demand response. Flexible operation enables aluminum smelters to concentrate electricity consumption during periods of high renewable energy generation, thereby improving the efficiency of renewable energy utilization and reducing the required renewable capacity to meet electricity demand. Second, from a reliability perspective, flexible operation also reduces the need for firm energy (energy from non-renewable resources that can be dispatched whenever needed [sepulveda_role_2018]) to maintain reliability and meet demand during periods of renewable scarcity. The reduction in non-renewable operational costs reflects this second mechanism, which reduces the need for coal- and/or natural gas-based generation (with carbon capture and storage) and nuclear power (see Supplementary 3.4 for details). However, non-renewable investment costs remain relatively unchanged, suggesting that flexible aluminum smelting primarily reduces the utilization of non-renewable resources, rather than replacing the need for non-renewable capacity.
@@
Comparing the different scenarios, we find that the flexibility value of aluminum smelting is higher in scenarios with high aluminum demand, because the aluminum smelting sector is bigger and the impact of load flexibility is therefore larger. As smelter operational flexibility increases, the cost savings from reduced battery capacity become larger; however, these are secondary to reductions in renewable investment and non-renewable operating costs. The overall electricity system cost savings are also similar between the Non-constrained-flexibility scenario and the Mid-flexibility scenario, suggesting diminishing returns in increasing smelter operational flexibility.
We also observe that flexible operation of aluminum smelting results in more investment in heating electrification (purple), such as the deployment of electric heat pumps. The increased investment reflects the fact that aluminum smelters can reduce winter peak loads through flexible operation, thereby enabling the greater deployment of clean heating technologies.
2.3 Cost-benefit analysis of retaining overcapacity

### v2

@@ Abstract
In many countries, declining demand in energy-intensive industries (EIIs) such as cement, steel, and aluminum is leading to industrial overcapacity. Although industrial overcapacity is traditionally envisioned as problematic and resource-wasteful, it could unlock EIIs’ flexibility in electricity use. Here, using China’s aluminum smelting industry as a case study, we evaluate the system-level cost-benefit of retaining EII overcapacity for flexible electricity use in decarbonized energy systems. We find that overcapacity can enable aluminum smelters to adopt a seasonal operation paradigm, ceasing production during winter load peaks that are exacerbated by heating electrification and renewable seasonality. This seasonal operation paradigm could reduce the investment and operational costs of China’s decarbonized electricity system by 23-32 billion CNY/year (11-15% of the aluminum smelting industry’s product value), sufficient to offset the increased smelter maintenance and product storage costs associated with overcapacity. It may also provide an opportunity for seasonally complementary labor deployment across the aluminum smelting and thermal power generation sectors, offering a potential pathway for mitigating socio-economic disruptions caused by industrial restructuring and energy decarbonization.
@@ 1 Introduction
Industrial restructuring — the transition of economies from heavy industry to higher-added-value sectors [54] — is leading to declining demand for industrial products and overcapacity in energy-intensive industries in many countries. China’s steel and cement industries, for instance, have averaged 20-30% overcapacity in the last decade, resulting in declining profitability and asset utilization [59]. At the same time, the deep decarbonization of these so-called “hard-to-abate sectors” requires structural shifts (e.g., increased recycling and electrification-driven technology substitution) that will exacerbate the overcapacity of existing primary (ore-based) production facilities [40, 38, 3, 57]. Studies project massive overcapacity for China’s aluminum smelters as recycled aluminum increasingly dominates supply (potentially fulfilling 70% of future demand [33, 34]). Traditionally, overcapacity is viewed as problematic, causing resource waste and economic risks [63]. Consequently, governments typically strive to eliminate overcapacity through policy measures [19], such as China’s Supply-Side Structural Reform (SSSR), which was initiated around 2016 to tackle massive industrial overcapacity.
@@
On the other hand, excess capacity also unlocks operational flexibility for energy-intensive industries, as it frees them from operating at full capacity at all times to meet demand. In decarbonized power systems reliant on variable renewable energy, which are inherently prone to multi-timescale (e.g., both diurnal and seasonal) supply-demand mismatches [9, 14], the demand-side flexibility of energy-intensive industries may become more valuable for balancing electricity supply and demand [27]. By shifting electricity consumption to periods of high renewable energy availability, energy-intensive industries can reduce both the energy system operational costs and carbon emissions [18]. Furthermore, unlike “fast-burst” demand-side resources such as electric vehicles and air conditioning [56], industrial facilities with overcapacity and material storage could potentially provide long-duration flexibility by scheduling their production over longer time horizons, functioning as a demand-side alternative to firm power or long-duration energy storage [45, 26]. In decarbonized energy systems with increasing variability in electricity prices [5, 12], this flexibility can also lower the energy input costs of energy-intensive industries like aluminum smelting, in which electricity costs exceed 30% of product costs [2].
To quantitatively assess this trade-off, we select China’s aluminum smelting industry as a representative case, because of its strong potential to become a source of flexibility for the power grid. Aluminum smelting is ten times more electricity-intensive per tonne of product than steelmaking [24, 20], and is also likely to have future overcapacity (over 50% by 2050 [33, 50, 16, 64]) that allows it to leverage its technical potential for operational flexibility. The aluminum industry is also becoming more integrated with an electricity grid that is increasingly dominated by renewables. Aluminum smelters have historically relied on on-site coal power and were located in northern coal hubs like Shandong [39]. However, they have recently been relocating to southern provinces like Yunnan, in order to take advantage of low-cost, renewable electricity in these regional grids [44]. Because of their massive electricity footprint and technical flexibility [55], these smelters have already been routinely targeted for mandated winter power curtailments during seasonal hydropower droughts [1]. Given these recent trends, aluminum smelting is an ideal archetype for examining the system-level benefits of overcapacity-enabled industrial flexibility.
Traditionally, demand-side flexibility has been viewed primarily as the ability to curtail or shift load for a few hours during extreme price spikes. However, deeply-decarbonized energy systems will also require long-term flexibility due to seasonal mismatch between electricity supply and demand. While extensive research has focused on supply-side seasonal balancing (e.g., long-duration thermal storage and hydrogen), long-duration flexibility on the demand side has received little attention. In the case of aluminum smelting, for instance, existing studies have primarily focused on validating the technical feasibility and potline-level thermal constraints of hourly-level flexible operation for short-term demand response [55, 58, 52, 15]. To the best of our knowledge, the feasibility and system-level value of long-term flexibility in smelting electricity use remain largely unexplored in current literature.
In this paper, we aim to understand the value of aluminum smelting overcapacity for providing demand flexibility in China’s decarbonized energy system. We examine these effects across both physical infrastructure and socio-economic dimensions (e.g., labor deployment), as both are essential for charting a sustainable path for the aluminum industry. However, the computational complexity of modeling and optimizing the intricate coupling between two already large and complex systems—the electricity system and aluminum smelters— is a significant barrier to such analysis and has not been done before. We developed an integrated provincial-level power-smelting co-optimization model (Figure 1). This approach couples the PyPSA power system framework [6] with a detailed aluminum smelter operation model [47] for hourly co-optimization. To solve this computationally heavy 8760-hour simulation, we adopted a Data-Driven Dimension Reduction (D3R) framework [37] paired with an iterative decomposition algorithm. First, D3R linearizes complex smelter constraints and generates an initial co-optimized solution to make the power system model easier to solve. The power system model then passes hourly locational marginal prices (LMPs) to the smelters. Using these price signals, each smelter independently optimizes its discrete production decisions (e.g., shutdown and restart costs) in parallel. This price-load feedback loop iterates until convergence, bridging modeling high-fidelity smelting physics with large-scale grid planning (see Section 4 for details).
Figure 1: Schematic diagram of the integrated power-aluminum co-optimization framework. To overcome the computational intractability of an 8760-hour horizon, the PyPSA-based grid model and detailed smelter operations are decoupled and solved iteratively through the exchange of price signals (LMPs) and industrial load profiles, aided by the D3R framework.
@@
We find that in decarbonized electricity systems, overcapacity enables aluminum smelters to shift operations away from periods of winter electricity supply scarcity, which are driven by low renewable availability and heating electrification. This flexibility lowers overall electricity system costs by reducing peak loads, which in turn decreases the generation capacity required to meet those peaks, as well as lowering utilization of coal, natural gas, and nuclear generation. In the Decarbonized Power 2050 (DP-2050) scenario, retaining 30% aluminum-smelting overcapacity reduces electricity system costs by 23-32 billion CNY/year (or 11-15% of the aluminum smelting industry’s product value), as compared to a no-overcapacity case. We find that, in such a scenario, aluminum production costs are lower by over 1,500 CNY/tonne (or 9% of the total production cost), due to aluminum smelters consuming cheaper electricity. We also explore the potential socio-economic implications of this transition. Because thermal power plants may also operate seasonally in decarbonized systems, complementary seasonal operation in the aluminum smelting industry reveals a potential for inter-industry labor deployment. Our analysis indicates that if workers could switch between industries on a seasonal basis, aggregated workforce volatility across the aluminum smelting and power generation sectors could theoretically be reduced by up to 25%. This highlights a possible co-benefit of seasonal aluminum smelter operation for bolstering social stability in deeply-decarbonized energy systems.
@@
Although we are using aluminum smelting as a case, our findings can be generalized to other energy-intensive industries, such as steel and cement industries, where seasonal operation could potentially bring significant energy input cost savings to offset the costs of maintaining overcapacity and product storage. More broadly, our work suggests that industrial overcapacity could be reframed from a problem to an opportunity to reduce energy system costs, lower industrial production costs, foster workforce stability, and thereby mitigate socio-economic disruptions from energy decarbonization.
@@ 2 Results
Smelter operational flexibility
Case Min. power consumption (per unit of nameplate value) Restart costs (EUR/MW)
Low $0.99$ 96,594
Mid $0.9$ 13,981
High $0.7$ 2,796
Unconstrained $0.0$ 0
Technology CAPEX
[…]
Low 624/346 186 520 575
Mid 780/432 232 650 719
High 780/432 348 975 1079
Labor and employment flexibility
Case Labor and standby costs Workforce mobility
Inflexible Constant regardless of output None (standby in smelters)
Flexible Reduced during shutdown Seasonal transfer to other sectors
The restart-up costs are measured from the electricity input side. Exchange rates throughout the study (annual averages 2020): 1 USD = 6.900 CNY; 1 EUR = 7.868 CNY; 1 USD = 0.877 EUR. The electricity system cost data is primarily from the PyPSA-China model [65], with the specific Chinese cost assumptions and their original sources documented in Supplementary Note 4.4. Long-duration storage includes H2 and pumped hydro storage. Our core scenario is defined by the combination of Mid primary aluminum demand, Mid smelter operational flexibility, Mid technology costs, and Inflexible labor.
This study employs a scenario analysis approach to explore the impact of key uncertainties on the system-level value of aluminum smelting overcapacity. As detailed in Table 1, we evaluate four primary dimensions: (i) primary aluminum demand, which determines the industry’s aggregate flexibility potential by defining the size of the sector; (ii) smelter operational flexibility, representing the physical minimum load constraints and the economic penalty (e.g., restart costs) of non-baseload operation; (iii) technology costs, reflecting the relative competitiveness of demand-side flexibility against supply-side flexibility alternatives (e.g., batteries and hydrogen); and (iv) labor flexibility, which dictates whether workforce costs can be mitigated during seasonal shutdowns. (see Method 4.1 for more details) To ensure the robustness and traceability of these scenario settings, the bounds for each dimension are systematically derived from empirical literature and industry standards. For technology costs, the “Mid” scenario adopts peer-reviewed trajectories from the PyPSA-China[65] dataset, while the “Low” and “High” bounds for the costs are scaled by 20-20% and +50+50%, respectively, adhering to the AACE International Class 4 estimate guidelines [4].
[…]
@@ 2.1 Inevitable overcapacity in aluminum smelting in the recycle era
Figure 2: China’s projected aluminum demand and smelting overcapacity under different scenarios. The top solid line represents total aluminum demand, the green area shows recycled aluminum production (RA), and the yellow area indicates primary aluminum production (PA) up to 2050. China’s aluminum smelting capacity at the end of 2024 (45 million tonnes (Mt) per year) is represented by the red dashed line. The area between the red dashed line (smelter capacity at 2024 level) and the yellow area represents the excess capacity of aluminum smelters unless prematurely decommissioned. Because smelters typically operate for over 50 years, the vast majority of China’s existing capacity will not reach its natural retirement age before 2050 (see Supplementary Note 3.3), effectively locking in this structural baseline. The overcapacity rate is represented by the purple line, and is given by one minus the utilization rate (1 - demand/capacity). Different scenarios are constructed based on various assumptions regarding economic development and aluminum product import/export levels, with the variable settings outlined in Table 1. Even under the high-demand scenario, China’s aluminum industry would retain 46% excess capacity — and as much as 86% under the low-demand scenario — posing a significant challenge to the sector.
@@
Our analysis shows that over the next one to two decades, China’s primary aluminum demand will rapidly decline and remain low as more aluminum is produced from scrap, leading to unprecedented overcapacity in the aluminum smelting industry. As shown in Figure 2, the proportion of recycled (secondary) aluminum production in total aluminum demand will increase from 25% in 2024 [8] to 60% by 2050. The increased recycled aluminum production stems from a significant increase in available scrap aluminum, which is a delayed effect of China’s climbing aluminum production over the past two decades, as aluminum products typically have a lifespan of 10-20 years [21]. While this transition entails substantial reductions in the aluminum industry’s energy consumption and carbon emissions, due to the lower energy intensity of producing aluminum from scrap [32], it would result in over 60% excess capacity for existing aluminum smelters. Furthermore, because aluminum smelters typically operate for over 50 years, the vast majority of China’s existing smelting capacity will be far from the retirement age, locking in this severe overcapacity (see Supplementary Note 3.3 for the detailed capacity age distribution).
2.2 Electricity system cost reduction compared to no-overcapacity case
@@
Figure 3: Reduction in electricity system costs due to overcapacity-enabled flexibility in aluminum smelting, compared to the no-overcapacity case. Colors represent different sources of changes in system cost: non-renewable investment costs (gray) and non-renewable operational costs (orange) account for the largest system cost reductions, reflecting a decreased reliance on firm generation resources (coal, gas, biomass, and nuclear) to ensure power supply reliability. In Mid-demand scenarios, cost reductions range from 23 to 32 billion CNY (11-15% of the aluminum smelting industry’s product value in 2050), across different smelter flexibility and technology cost settings. Conversely, the observed increases in renewable energy and heating electrification investments indicate that seasonal load-shifting makes these clean technologies more economically efficient. By concentrating electricity consumption during summer periods of abundant renewable generation and avoiding winter peak loads, smelter flexibility eliminates the need for expensive firm generation and massive storage. This allows the system to cost-effectively expand renewable capacity and clean heating without facing severe integration barriers.
@@
Figure 3 shows the reduction in electricity system costs in the DP-2050 scenario when the operation of aluminum smelters with overcapacity is co-optimized with the electric power system, as compared to the no-overcapacity case. This represents the value of flexibility in smelting electricity use. Here, the power system cost includes operational costs and the annualized capital investment for new generation, storage, and transmission capacity. The costs of the aluminum smelters are not included in the electricity system costs, but are discussed later in the paper. The results show that in the Mid-demand scenario, the overcapacity of aluminum smelters provides a flexibility value of 23 to 32 billion CNY to the power system. To put this number into perspective, this is equivalent to 11-15% of the aluminum smelting industry’s product value in 2050.
Across all scenarios, the cost reductions are driven by substantial decreases in both the investment and operational costs (mainly fuel costs) of non-renewable energy sources (coal, natural gas, and nuclear). This is accompanied by increased investments in renewable energy and heating electrification in order to reach decarbonization constraints. Two interconnected mechanisms explain this shift toward a more renewable-dominated portfolio. First, without adequate long-duration demand-side flexibility, the expansion of variable wind and solar power faces diminishing marginal benefits, as their high variability and mismatch with electricity consumption lead to severe curtailment or the need for massive storage. By concentrating electricity consumption during periods of high renewable generation in the summer, flexible smelter operation enhances the utilization efficiency of renewables, making it economically beneficial to deploy greater renewable capacity. Second, shifting smelting load to the summer mitigates winter peak loads, significantly reducing the need for costly clean-firm energy (dispatchable low-carbon resources such as coal or natural gas with CCS, biomass with CCS, or fossil power offset by direct air capture [45]) during periods of renewable scarcity.
@@
Comparing the different scenarios, we find that the flexibility value of aluminum smelting is higher in scenarios with high aluminum demand, because the aluminum smelting sector is bigger and the impact of load flexibility is therefore larger. The overall electricity system cost savings are similar between the Non-constrained-flexibility scenario and the Mid-flexibility scenario, suggesting diminishing returns in increasing smelter operational flexibility.
2.3 Cost-benefit analysis of retaining smelting overcapacity